# Visual Agents as Fast and Slow Thinkers

### v2

@@ 5 Conclusion
In this study, we introduced FaST, an innovative framework that integrates the concepts of System 1 (fast, intuitive) and System 2 (slow, deliberate) thinking into visual agents, aiming to enhance their reasoning and decision-making capabilities. FaST dynamically adapts to queries of varying complexity by employing a flexible system switch mechanism, ensuring rapid responses for straightforward questions and employing hierarchical reasoning for more intricate scenarios. Our results across a wide array of benchmarks demonstrate significant improvements in performance, highlighting the efficacy of FaST’s chain of evidence methodology in reducing hallucination and enhancing interpretability. Our ablation studies highlight the crucial role of contextual clues and pixel-level adapters in achieving comprehensive visual reasoning and understanding. FaST significantly advances visual agents’ capability, paving the way for the more reliable, accurate, and human-like AI cognition process.
@@ Appendix A Implementation Details and Pseudo-code of FaST
Hyper-parameters. We follow established methodologies [26] and utilize LLaVA-v1.5  [26] as the foundational visual agent. The image resolution is preprocessed to 336 × 336 pixels to accommodate the clip-vit-large-patch14-336 vision encoder [98]. The AdamW optimizer [120] is employed with the DeepSpeed ZeRO 2  [^1] https://github.com/microsoft/DeepSpeed configuration for fine-tuning the switch, proposal, and segmentation adapters with LoRA [113]. For the LoRA configuration, we set the rank to 128 and alpha to 256, consistent with the settings of LLaVA-v1.5. Additionally, we adjust the learning rate of the vision encoder projection layer to 2e-5 to achieve better alignment. An MLP projection with channels of [256, 4096, 4096] is used to connect image representations into the word embedding space for the projection layer. An additional resampler projection layer is used to reduce the number of image tokens.
@@
Training Data for Switch Adapter. Consistent with the pretraining stage of LLaVA-v1.5, we initially pretrain Vicuna-v1.5 as a base frozen large language model and for the MLP projection layer and sampler layer of the CLIP vision encoder using a 558K subset of the LAION-CC-SBU dataset  [^2] https://huggingface.co/datasets/liuhaotian/LLaVA-Pretrain with BLIP [121] captions. During the fine-tuning stage, we integrate the negative dataset acquired from $V^{*}$ [19] with the original LLaVA-v1.5 instruction tuning 665k data [^3] https://huggingface.co/datasets/liuhaotian/LLaVA-Instruct-150K for LoRA based finetuning.
@@
Training Data for Proposal Adapter To determine the corresponding region for a query, we use LRP++ [122] for data construction, similar to Chain of Spot [123]. Our initial prompt is as follows:
@@
Training Data for Seg Adapter. Adopting an approach similar to LISA [29], the training data for our model comprises three distinct segments: a semantic segmentation dataset, a referring segmentation dataset, and a reasoning segmentation dataset. We deliberately exclude visual question-answering datasets to enhance the model’s segmentation performance. The semantic segmentation segment includes the ADE20K [100], COCO-Stuff [99], and LVIS-PACO [101] part segmentation datasets. The referring segmentation datasets encompass refCOCO[105], refCOCO+ [105], refCOCOg [99], and refCLEF [124]. The reasoning segmentation dataset includes ReasonSeg [29]. It is important to note that the referring segmentation and reasoning segmentation datasets are carefully excluded during the evaluation of the segmentation benchmarks to prevent any potential data leakage.
@@
Appendix B More Qualitative Results for Visual Question Answering
@@
Figure 8: Failure cases of FaST.
Appendix C More Qualitative Results for Segmentation
@@
Appendix D Failure Case
In Fig. 8, we present an overview of the most notable failure cases, providing insights into the distinct patterns that lead to suboptimal outputs in our FaST model. These challenges include difficulty in triggering the System 2 thinking mode, constructing adequate contextual clues, generating appropriate proposals, and providing accurate pixel masks. The model often fails to recognize the need for deliberate reasoning, relying instead on System 1 thinking, which leads to incorrect responses, as seen in Fig. 8. Inadequate contextual clues generated by the switch adapter impair the model’s focus on the correct region, resulting in vague or incorrect responses, as illustrated in Fig. 8. The proposal adapter’s inaccurate identification of regions of interest, as shown in Fig. 8, leads to proposals that do not correspond to the query. Additionally, the segmentation adapter struggles with producing precise masks, particularly for small or occluded objects, causing erroneous conclusions, as highlighted in Fig. 8. These failure cases underscore the urgent need for refinement in our FaST framework, emphasizing the importance of significantly enhancing the precision of the system switch adapter, improving contextual clue construction, and optimizing the proposal and segmentation adapters to achieve more reliable and consistently accurate responses in complex visual and textual tasks.
Appendix E Limitation and Future Work
@@
Appendix F Social Impacts
The development and deployment of FaST carry significant implications for the field of AI, especially in augmenting the human-like cognitive capabilities of LLM-based visual agents. On the positive side, FaST’s emulation of human cognitive processes holds promise for creating more advanced AI systems that can assist in critical areas like general vision question dialogue and security surveillance, where precise visual comprehension is paramount. Moreover, its transparent decision-making pipeline contributes to the explainability of AI, fostering trust and facilitating ethical AI practices. However, it’s imperative also to acknowledge potential negative social impacts. The reliance on extensive models and datasets could perpetuate biases if not meticulously curated, potentially resulting in unjust or discriminatory outcomes. Addressing these ethical considerations and establishing guidelines for the responsible utilization of such sophisticated AI systems are paramount requirements.
Appendix G Ethical Safeguards
@@
Appendix H Reproducibility
Our FaST framework is implemented in PyTorch [125]. All the experiments are conducted on eight NVIDIA A100-80GB GPUs. Our full implementation shall be publicly released upon paper acceptance to guarantee reproducibility. The codes are available at the anonymous link https://anonymous.4open.science/r/Sys2-LLaVA-8B0F/ for the review process.
@@ Appendix H Reproducibility
Appendix I Licenses for existing assets

### v3

@@ 5 Conclusion
In this study, we introduced FaST, a framework that combines System 1 (which is fast and intuitive) and System 2 (which is slow and deliberate) thinking to improve visual agents’ reasoning and decision-making. FaST adapts to queries of varying complexity with a flexible system switch, delivering quick responses for simple tasks and using hierarchical reasoning for more complex scenarios. The FaST leverages neuro-symbolic decision-making transparent pipeline delivering interpretable intermediate outputs that enable explainability. Our results show significant improvements across benchmarks, demonstrating the effectiveness of FaST’s chain of evidence in reducing hallucinations and improving interpretability. Furthermore, ablation studies highlight the critical importance of contextual clues, symbolic reasoning, and pixel-level adapters in refining visual reasoning and understanding, marking a step forward in creating more reliable and accurate AI cognition.
6 Ethical Safeguards
7 Reproducibility
Our FaST framework is implemented in PyTorch (Paszke et al., 2019). All the experiments are conducted on eight NVIDIA A100-80GB GPUs. Our full implementation shall be publicly released upon paper acceptance to guarantee reproducibility. The codes are available at the anonymous link https://anonymous.4open.science/r/Sys2-LLaVA-8B0F/ for the review process.
All Experiments are conducted on eight NVIDIA A100-80GB SXM GPUs [^1] https://www.nvidia.com/en-sg/data-center/a100/. Reproducing the fine-tuning process would require approximately 15 A100 GPU days.
@@ Appendix A Implementation Details and Pseudo-code of FaST
Visual Resampler. The resampler (Alayrac et al., 2022) compresses high-dimensional visual features into a fixed-size latent space using a cross-attention mechanism. It begins with a set of learnable latent embeddings, which query the vision encoder’s output features through scaled dot-product attention. Each latent embedding attends selectively to the most relevant visual tokens, guided by attention weights computed via the query-key interaction. The process iterates across multiple layers of cross-attention, followed by feedforward transformations, refining the latent representations at each step. This approach ensures efficient dimensionality reduction while retaining critical information, producing a compact set of visual tokens for downstream tasks.
@@
Hyper-parameters. We follow established methodologies and utilize LLaVA-v1.5  (Liu et al., 2024a) as the foundational visual agent. The image resolution is preprocessed to 336 × 336 pixels to accommodate the clip-vit-large-patch14-336 vision encoder (Radford et al., 2021). The AdamW optimizer (Loshchilov & Hutter, 2019) is employed with the DeepSpeed ZeRO 2  [^2] https://github.com/microsoft/DeepSpeed configuration for fine-tuning the switch, proposal, and segmentation adapters with LoRA (Hu et al., 2022). For the LoRA configuration, we set the rank to 128 and alpha to 256, consistent with the settings of LLaVA-v1.5. Additionally, we adjust the learning rate of the vision encoder projection layer to 2e-5 to achieve better alignment. An MLP projection with channels of [256, 4096, 4096] is used to connect image representations into the word embedding space for the projection layer. An additional resampler projection layer is used to reduce the number of image tokens.
@@
Training Data for Switch Adapter. Consistent with the pretraining stage of LLaVA-v1.5, we initially pretrain Vicuna-v1.5 as a base frozen large language model and for the MLP projection layer and sampler layer of the CLIP vision encoder using a 558K subset of the LAION-CC-SBU dataset  [^3] https://huggingface.co/datasets/liuhaotian/LLaVA-Pretrain with BLIP (Li et al., 2022) captions. During the fine-tuning stage, we integrate the negative dataset acquired from $V^{*}$ (Wu & Xie, 2024) and PixelLM (Ren et al., 2024) and with the original LLaVA-v1.5 instruction tuning 665k data [^4] https://huggingface.co/datasets/liuhaotian/LLaVA-Instruct-150K for LoRA based finetuning.
Specifically, The dataset for fine-tuning the switch adapter was carefully constructed to emphasize scenarios requiring precise object recognition and complex reasoning. For the GQA subset of the 167k VQA data, we specifically targeted questions where the annotated objects mentioned in the query were critical for deriving the correct answer. Initially, the InstructBLIP model was used to evaluate GQA questions with annotated objects. Only questions that the model could correctly answer were retained. To ensure the importance of these annotated objects, we applied the LaMa image inpainting model to erase the mentioned objects from the corresponding images. The modified images were re-evaluated using InstructBLIP, and only questions that the model failed to answer after object removal were included. This process ensured that the curated subset focused exclusively on questions where the annotated objects were essential, forming a robust component of the VQA data.
For the VAW object attribution dataset, both open-ended and binary questions about object attributes were synthesized. Open-ended questions were formulated around attributes such as “color,” “material,” and “pose,” while binary questions incorporated additional attributes like “state” and “optical property.” Answer formats adhered to predefined structures to ensure consistency. The same object removal and re-evaluation strategy as used in the GQA subset was applied, filtering the data to include only questions where the absence of objects rendered the query unanswerable.
From the LLaVA-80K instruction tuning data, noun phrases were extracted from the text of questions or instructions and matched with object category names defined by COCO, augmented with common synonyms such as “man” and “woman” for the “person” category. Images were retained only if the identified categories had annotated instances with bounding boxes. These annotated instances, along with their bounding box coordinates, were used as target objects during training.
In addition, we incorporated a data generation approach inspired by LISA, utilizing GPT-4 and GPT-4V to expand and diversify the dataset. Initially, LLAVA was used for image captioning, and GPT-4 generated questions about multiple regions in the image. While this approach utilized pre-existing mask annotations to reduce costs, its diversity was limited to the scope of the captions. To address these limitations, we refined the pipeline with GPT-4V, leveraging its advanced capabilities in visual understanding. Image captions, object names, and bounding box coordinates were input into GPT-4V, which, using dynamically crafted prompts, autonomously selected instances and generated nuanced question-answer pairs tailored to the image content. This refinement significantly improved the diversity and contextual relevance of the data. An illustrative example of such prompts is provided below:
Prompt: Imagine you need to query a machine agent about an image. The image has a height of 720 pixels and a width of 1280 pixels. You are given several entities described by a list, each identifying an object in the image along with its location. The class names and corresponding coordinates are as follows:
• Dog at [350.12, 450.45, 480.89, 600.67];
• Ball at [200.33, 300.22, 250.78, 350.56];
• Grass at [0.0, 500.0, 1280.0, 720.0];
Coordinates are represented as (top-left x, top-left y, bottom-right x, bottom-right y). The question must incorporate at least two of these objects and require reasoning about the relationships or interactions between them. Additional requirements for the generated question are as follows:
1.The answer to the question must explicitly reference each included object or its equivalent and avoid implying the presence of any other objects not listed.
2.The question must be precise, meaningful, and avoid being overly general.
3.The question should frame a single cohesive activity or relationship rather than merely combining independent sub-queries.
4.When answering, the class names should be rephrased to indicate their position, role, or interaction in the image.
This multi-faceted dataset construction process ensured the generation of diverse and challenging samples, providing a robust foundation for fine-tuning the switch adapter on complex reasoning tasks.
@@
Training Data for Proposal Adapter To determine the corresponding region for a query, we use LRP++ (Chefer et al., 2021) for data construction, similar to Chain of Spot (Liu et al., 2024b). Our initial prompt is as follows:
@@
Training Data for Seg Adapter. Adopting an approach similar to LISA Lai et al. (2024), the training data for our model comprises three distinct segments: a semantic segmentation dataset, a referring segmentation dataset, and a reasoning segmentation dataset. We deliberately exclude visual question-answering datasets to enhance the model’s segmentation performance. The semantic segmentation segment includes the ADE20K (Zhou et al., 2017), COCO-Stuff (Caesar et al., 2018), and LVIS-PACO (Ramanathan et al., 2023) part segmentation datasets. The referring segmentation datasets encompass refCOCOKazemzadeh et al. (2014), refCOCO+ (Kazemzadeh et al., 2014), refCOCOg (Caesar et al., 2018), and refCLEF (Rohrbach et al., 2016). The reasoning segmentation dataset includes ReasonSeg (Lai et al., 2024). It is important to note that the referring segmentation and reasoning segmentation datasets are carefully excluded during the evaluation of the segmentation benchmarks to prevent any potential data leakage.
@@
Appendix B More Results for Different Thinking Modes
As shown in Table 6, FaST demonstrates strong performance in both System 1 and System 2 reasoning on VQA datasets, outperforming baseline methods in most cases. Notably, FaST achieves the highest accuracy in challenging System 2 tasks across GQA, $VQA^{T}$, and $SQA^{I}$, which require advanced reasoning capabilities. This highlights the effectiveness of the switch adapter mechanism in dynamically allocating tasks based on complexity. While maintaining competitive performance in simpler System 1 tasks, FaST leverages its adaptive architecture to excel in more complex scenarios, as evidenced by its superior System 2 results.
Further, in Table 7, FaST’s robustness extends to reasoning segmentation tasks, where it achieves significant improvements in System 2 performance compared to baseline models such as LLaVA with segmentation and LISA-7B. For example, in the ReasonSeg dataset, FaST records a remarkable 48.2 CIoU in System 2 tasks, significantly outperforming LISA-7B and LLaVA, which achieve 43.3 CIoU and 42.4 CIoU, respectively. This result underscores FaST’s ability to generalize effectively across diverse task families and reasoning paradigms.
Overall, the results validate the universal applicability and robustness of the FaST framework. By effectively utilizing the switch adapter to allocate tasks dynamically, FaST demonstrates a strong capability to balance performance across both simple and complex reasoning tasks, making it a reliable solution for diverse real-world applications.
  $VQA^{v2}$ GQA $VQA^{T}$ $SQA^{I}$ Method LLM Sys 1 Sys 2 Sys 1 Sys 2 Sys 1 Sys 2 Sys 1 Sys 2 BLIP-2 Vicuna-13B 67.3 53.1 37.8 22.4 44.3 39.7 63.4 59.2 LLaVA-v1.5 Vicuna-7B 81.2 68.0 70.3 47.0 61.1 53.7 68.4 65.7 Chain of Spot Vicuna-7B 82.1 74.5 70.9 50.7 62.1 59.0 68.6 67.8 FaST (Ours) Vicuna-7B 81.1 75.5 70.2 52.3 61.2 60.2 68.2 70.2
Table 6: System 1 and System 2 performance on VQA datasets. FaST demonstrates superior performance in both reasoning modes compared to baselines.
  refCOCOg ReasonSeg Method Sys 1 Sys 2 Sys 1 Sys 2 LISA-7B 70.2 63.4 46.6 43.3 LLaVA w Seg 68.4 60.2 44.2 42.4 FaST (Ours) 70.8 64.1 46.4 48.2
Table 7: System 1 and System 2 performance on reasoning segmentation tasks. FaST achieves strong performance across both tasks, demonstrating its robustness and effectiveness in dynamic task allocation between System 1 and System 2.
Appendix C More Qualitative Results for Visual Question Answering
@@
Figure 9: Failure cases of Our FaST system.
Appendix D More Qualitative Results for Segmentation
@@
Appendix E Failure Case
In Figure 9, we present an overview of the most notable failure cases, providing insights into the distinct patterns that lead to suboptimal outputs in our FaST model. These challenges include difficulty in triggering the System 2 thinking mode, constructing adequate contextual clues, generating appropriate proposals, and providing accurate pixel masks. The model often fails to recognize the need for deliberate reasoning, relying instead on System 1 thinking, which leads to incorrect responses, as seen in Figure 9. Inadequate contextual clues generated by the switch adapter impair the model’s focus on the correct region, resulting in vague or incorrect responses, as illustrated in Fig. 9. The proposal adapter’s inaccurate identification of regions of interest, as shown in Figure 9, leads to proposals that do not correspond to the query. Additionally, the segmentation adapter struggles with producing precise masks, particularly for small or occluded objects, causing erroneous conclusions, as highlighted in Figure 9. These failure cases underscore the urgent need for refinement in our FaST framework, emphasizing the importance of significantly enhancing the precision of the system switch adapter, improving contextual clue construction, and optimizing the proposal and segmentation adapters to achieve more reliable and consistently accurate responses in complex visual and textual tasks.
Appendix F Limitation and Future Work
@@
For the recent large reasoning model like OpenAI o1 model, while these models leverage reinforcement learning and internal chains of thought to achieve scalability, they often require significant computational resources, making them less efficient.
In contrast, FaST is designed to prioritize multimodal reasoning with a clear focus on transparency and adaptability. The use of interpretable reasoning modes (System 1 and System 2) ensures that FaST provides insights into its decision-making processes, which is critical for applications requiring explainability. Additionally, FaST’s modular design allows it to balance computational efficiency and accuracy dynamically, making it suitable for diverse and resource-constrained environments. These strengths highlight FaST’s unique contributions and its complementary potential to scalable reasoning approaches.
Appendix G Social Impacts
The development and deployment of FaST significantly advance AI by enhancing the human-like cognitive abilities of LLM-based visual agents. Positively, FaST enables sophisticated applications in areas like vision-based dialogues and security surveillance, while its transparent decision-making fosters trust and ethical AI practices. However, reliance on large models and extensive datasets risks perpetuating biases, potentially leading to unjust or discriminatory outcomes. Addressing these ethical concerns and establishing responsible usage guidelines are essential for the responsible deployment of such advanced AI systems.
Appendix H Ethical Safeguards
@@
Appendix I Reproducibility
Our FaST framework is implemented in PyTorch (Paszke et al., 2019). All the experiments are conducted on eight NVIDIA A100-80GB GPUs. Our full implementation shall be publicly released upon paper acceptance to guarantee reproducibility. The codes are available at the anonymous link https://anonymous.4open.science/r/Sys2-LLaVA-8B0F/ for the review process.
@@ Appendix I Reproducibility
Appendix J Licenses for existing assets